# LP-ICP: General Localizability-Aware Point Cloud Registration for Robust Localization in Extreme Unstructured Environments

### v2

@@ III-C System Overview
The Optimization-Module (Section V) utilizes the categorization results from the LocDetect-Module for each direction of the 6-DOF pose, along with higher localizability contributions. It constructs optimization equations to solve for the optimal pose estimate $\boldsymbol{x}^{*}$. We use a method combining soft and hard constraints. Hard constraints are applied in non-localizable directions to prevent pose updates in those directions. Soft constraints are added in partially localizable directions, allowing the pose to be updated under the constraint, with the pose tending toward the constraint value rather than being strictly equal to it. This reduces the risk of large pose fluctuations or drift, which could otherwise lead to localization failure. The pose update in fully localizable directions is unaffected.
@@ IV-A Localizability of an Individual Correspondence Constraint
Since $\boldsymbol{H}$ is a real symmetric matrix, it can be diagonalized by eigenvalue decomposition as follows:
@@
$\boldsymbol{\Lambda}$ $=(\boldsymbol{J}\boldsymbol{V})^{T}\boldsymbol{J}\boldsymbol{V}$ (9)
@@
We consider (11) and (14), as well as (12) and (13), to be similar in nature but different in value. When (13) and (14) are used in the point-to-plane ICP algorithm, and rotation is represented using Lie algebra, they become identical to the localizability evaluation formula of X-ICP [21]. In other words, the localizability detection formula of [21] is a special case of the conclusions derived in this study.
For the convenience of subsequent calculations and usage, we choose to decouple translation and rotation in the above derivation. That is, for $N$ correspondence constraints, we have: $\boldsymbol{J}_{r}=(\boldsymbol{J}_{r1}^{T}\cdots\boldsymbol{J}_{rN}^{T})^{T}\in\mathbb{R}^{N\times 3}$ in the rotation directions and $\boldsymbol{J}_{t}=(\boldsymbol{J}_{t1}^{T}\cdots\boldsymbol{J}_{tN}^{T})^{T}\in\mathbb{R}^{N\times 3}$ in the translation directions, where $\boldsymbol{J}_{ri}\in\mathbb{R}^{1\times 3}$ is the Jacobian of the i-th residual $f_{i}\boldsymbol{(}x)$ w.r.t. the 3-DOF rotation and $\boldsymbol{J}_{ti}\in\mathbb{R}^{1\times 3}$ is the Jacobian of the i-th residual $f_{i}\boldsymbol{(}x)$ w.r.t. the 3-DOF translation. That is, it is equivalent to performing eigenvalue decomposition on the top-left 3$\times$3 matrix $\boldsymbol{H}_{r}$ and the bottom-right 3$\times$3 matrix $\boldsymbol{H}_{t}$ of the Hessian matrix $\boldsymbol{H}$, as follows:
@@
$\lambda_{j}=\sum_{i=1}^{N}(\boldsymbol{J}_{ri}\boldsymbol{v}_{rj})^{2}$ (17)
@@
$\lambda_{j}=\sum_{i=1}^{N}(\boldsymbol{J}_{ti}\boldsymbol{v}_{tj})^{2}$ (19)
@@ IV-B3 Localizability Contribution Vector
$\boldsymbol{F}_{r}=\left[\boldsymbol{F}_{r1}^{e}\cdots\boldsymbol{F}_{rN_{1}}^{e}\;\boldsymbol{F}_{r1}^{p}\cdots\boldsymbol{F}_{rN_{2}}^{p}\right]^{T}$ (31)
$\boldsymbol{F}_{t}=\left[\boldsymbol{F}_{t1}^{e}\cdots\boldsymbol{F}_{tN_{1}}^{e}\;\boldsymbol{F}_{t1}^{p}\cdots\boldsymbol{F}_{tN_{2}}^{p}\right]^{T}$ (32)
@@ IV-C1 Filtering Out Noise and Picking Out High Localizability Contributions
An example of the calculation results for $L_{f}(j)$ and $L_{u}(j)$ is provided in Fig. 3. They are computed from the localizability contributions of a LiDAR scan performed in an open area with slightly undulating terrain. In the directions along $\boldsymbol{v}_{r1}$, $\boldsymbol{v}_{r2}$ and $\boldsymbol{v}_{t1}$, the values of $L_{f}(j)$ and $L_{u}(j)$ are large, indicating well-constrained conditions. In the directions along $\boldsymbol{v}_{r3}$, $\boldsymbol{v}_{t2}$ and $\boldsymbol{v}_{t3}$, the values of $L_{f}(j)$ and $L_{u}(j)$ are small, indicating ill-conditioned constraints. In ill-conditioned directions, the localizability contribution from edge point-to-line correspondences (red region) dominates, especially in the calculation of $L_{u}(j)$, which represents the sum of high localizability contributions. Therefore, it is essential to utilize the localizability information from edge point-to-line correspondences.
Figure 3: Two exemplary double Y-axis histograms presenting $L_{f}$ and $L_{u}$ along each eigenvector direction of the 6-DOF pose. The calculated localizability contributions are from a LiDAR scan. It indicates smaller values of $L_{f}(j)$ and $L_{u}(j)$ in the $\boldsymbol{v}_{r3}$, $\boldsymbol{v}_{t2}$ and $\boldsymbol{v}_{t3}$ directions. The red region represents the localizability contribution from edge point-to-line correspondences, while the pink region represents the contribution from planar point-to-plane correspondences.
Figure 4: An example of the Localizability Detection Module. (a) Scan: A LiDAR scan obtained in an open area with slightly undulating terrain. (b) Localizability Contribution: Two double Y-axis histograms identical to those in Fig. 3. They are calculated based on the localizability contribution of the scan in (a). (c) Categorization: The localizability category along each eigenvector direction of the 6-DOF pose is assigned through conditional judgment. The threshold is set according to Section IV-C. (d)Result: The results of the localizability categories along the eigenvector directions are shown.
@@ IV-C2 Categorization
That is, when $L_{f}(j)\geq T_{1}\text{ or }L_{u}(j)\geq T_{2}$, the localizability category of the direction along $\boldsymbol{v}_{j}$ is considered fully localizable, i.e. the category is “Full”. When $L_{f}(j)\geq T_{3}\text{ and }L_{u}(j)\geq T_{4}$, the localizability category of the direction along $\boldsymbol{v}_{j}$ is considered partially localizable, i.e. the category is “Partial”. In this direction, both $L_{f}(j)$ and $L_{u}(j)$ must meet their respective thresholds. This means that there must be a sufficient number of points with high localizability contributions in the direction along $\boldsymbol{v}_{j}$, which helps improve the optimization quality after additional constraints are added in the subsequent steps. Otherwise, the localizability category of the direction of $\boldsymbol{v}_{j}$ is considered non-localizable, i.e. the category is “None”. In this study, based on experimental validation, for all subsequent experiments, $T_{1}$ is set to 50, $T_{2}$ is set to 30, $T_{3}$ is set to 15 and $T_{4}$ is set to 9.
@@
The categories for directions along $\boldsymbol{v}_{r1}$, $\boldsymbol{v}_{r2}$ and $\boldsymbol{v}_{t1}$ are fully localizable, indicating good geometric constraints. the categories for directions along $\boldsymbol{v}_{r3}$, $\boldsymbol{v}_{t2}$ are partially localizable, suggesting that although these directions exhibit degeneracy, some geometric information is still usable. The category for directions along $\boldsymbol{v}_{t3}$ is non-localizable, indicating severe degeneracy in this direction with minimal usable geometric information.
@@ V Optimization Module
The Optimization Module utilizes the localizability information provided by the LocDetect-Module. Based on the localizability categories of the 6-DOF pose along the directions of the eigenvectors, additional soft and hard constraints are incorporated into the optimization equation to achieve robust pose estimation. Specifically, in partially localizable directions, soft constraints are added using constraint values computed from higher localizability contributions. This allows the pose to trend toward the constraint values without being strictly equal, thereby avoiding significant pose fluctuations and drift that could result in localization failure. In non-localizable directions, hard constraints are applied to prevent pose updates. In fully localizable directions, pose updates remain unrestricted.
Note that to improve time efficiency in the subsequent experiments, the localizability category determination and additional constraint calculations are performed only during the first iteration of each ICP execution. For all subsequent iterations before convergence, the categories and additional constraint information computed in the first iteration are used directly in the optimization. Because in most SLAM algorithms, ICP-based pose estimation is typically performed at a relatively high frequency (e.g., 10 Hz in LVI-SAM), this configuration has little impact on accuracy.
@@ V-A Optimization Strategy
Zhang et al.[20] uses binary degeneracy judgment and applies solution remapping to set the pose update in degenerate directions to zero. However, it has the limitation of insufficient utilization of constraint information. [21] classifies localizability into three categories. During optimization, it adds hard constraints in partially localizable directions, forcing the update value to be equal with the pose update calculated from a small subset of high-localizability points, thus utilizing more constraint information. However, we find that when integrating X-ICP into some SLAM systems, this approach increases the risk of large fluctuations in pose updates, which can easily cause localization failure.
To address the above issue, we use an optimization method that combines soft and hard constraints in the Optimization-Module, utilizing the localizability information from the LocDetect-Module. Each time the ICP algorithm starts, the process begins with the initial pose estimate $\boldsymbol{x}_{0}$, which can be obtained through various methods in SLAM, such as previous odometry. Through iteration, the goal is to find the optimal pose update $\Delta\boldsymbol{x}$, which is added to $\boldsymbol{x}_{0}$ to obtain the optimal pose solution $\boldsymbol{x}^{*}$.
@@
    Partially Localizable (Partial): First, obtain the edge point-to-line correspondences and planar point-to-plane correspondences with moderate localizability contributions in the direction of $\boldsymbol{v}_{j}$ from the LocDetect-Module (i.e., satisfying (33)). Then, use them for a small local ICP optimization, which combines point-to-line and point-to-plane cost functions, to obtain an additional constraint value $\Delta\boldsymbol{r}_{0}\in\mathbb{R}^{3\times 1}$ or $\Delta\boldsymbol{t}_{0}\in\mathbb{R}^{3\times 1}$ for the pose update in the direction of $\boldsymbol{v}_{j}$. Then, a soft constraint is added with weights, making the final pose update along the $\boldsymbol{v}_{j}$ direction tend toward the constraint value. The specific process is described in Section V-B.
@@ V-B Adding Soft Constraints in Partially Localizable Directions
    A soft constraint is added to the final optimization equation, ensuring that the final pose update in the direction of $\boldsymbol{v}_{j}$ tends to approach the projection of the constraint value $\Delta\boldsymbol{r}_{0}$ or $\Delta\boldsymbol{t}_{0}$ in the direction of $\boldsymbol{v}_{j}$. A squared loss function is used, as follows:
@@
$\mu_{j}$ is the set weight. By adjusting the value of $\mu_{j}$, the extent to which the final pose update’s projection in the $\boldsymbol{v}_{j}$ direction tends toward the projection of the constraint values $\Delta\boldsymbol{r}_{0}$ or $\Delta\boldsymbol{t}_{0}$ in the $\boldsymbol{v}_{j}$ direction can be controlled. In this paper, $T_{5}$ is set to 15 for all subsequent experiments. When $L_{u}(j)$ in the direction of $\boldsymbol{v}_{j}$ is greater than $T_{5}$, it indicates that more correspondences with high localizability contributions are used to compute $\Delta\boldsymbol{r}_{0}$ or $\Delta\boldsymbol{t}_{0}$, making the calculated values more reliable, and thus the weight $\mu_{j}$ is increased. If the quality and quantity of the correspondences used in (38) or (39) are high, the weight $\mu_{j}$ can be appropriately increased. However, if $\mu_{j}$ is too large, it may cause fluctuations and jaggedness in the optimized trajectory, leading to instability. The setting of $\mu_{j}$ should ensure that the pose estimate in the direction of $\boldsymbol{v}_{j}$ tends toward the constraint value, while also maintaining the smoothness and stability of the trajectory.
@@ VI Experiment
In this section, the experimental setup is introduced in Section VI-A. The proposed method is then evaluated through experiments on a simulated planetary-like environment dataset (Section VI-B) and a real-world underground tunnel dataset (Section VI-C), with comparisons to state-of-the-art methods. Finally, the runtime performance of the algorithm is evaluated (Section VI-D).
@@ VI-A1 Implementation Details
The proposed framework, LP-ICP, is integrated into the point cloud registration framework of LVI-SAM [11]. LVI-SAM combines LIO and VIO sub-modules, with the LIO sub-module serving as the main component, while the VIO sub-module provides initial estimates for the LIO sub-module. The LIO module performs scan-to-map registration using edge point-to-line and planar point-to-plane correspondences. All experiments and evaluations in this paper are performed on a laptop equipped with an Intel i7-12700H CPU.

### v3

@@ III-C System Overview
The Optimization-Module (Section V) utilizes the categorization results from the LocDetect-Module for each direction of the 6-DOF pose, along with higher localizability contributions. It constructs optimization equations to solve for the optimal pose estimate $\boldsymbol{x}^{*}$. We use a method combining soft and hard constraints. Hard constraints [22] are applied in non-localizable directions to prevent pose updates in those directions. Soft constraints are added in partially localizable directions, allowing the pose to be updated under the constraint. The pose update in fully localizable directions is unaffected.
@@ IV-A Localizability of an Individual Correspondence Constraint
Since $\boldsymbol{H}$ in (7) is a real symmetric positive definite matrix, it can be diagonalized by eigenvalue decomposition as follows:
@@
$\boldsymbol{\Lambda}$ $=(\boldsymbol{J}\boldsymbol{V})^{T}\boldsymbol{J}\boldsymbol{V}$
@@
We consider (9) and (12), as well as (10) and (11), to be similar in nature but different in value. When (11) and (12) are used in the point-to-plane ICP algorithm, and rotation is represented using Lie algebra, they become identical to the localizability evaluation formula of X-ICP [22].
For the convenience of subsequent calculations and usage, we choose to decouple translation and rotation in the localizability analysis in the above derivation. Tuna et al. [22] provides an explanation for performing localizability analysis separately in the translational and rotational eigenspaces from the perspective of information analysis. Note that we do not explicitly handle cases where the eigenvectors couple rotational and translational states. For $N$ correspondence constraints, we have: $\boldsymbol{J}_{r}=(\boldsymbol{J}_{r1}^{T}\cdots\boldsymbol{J}_{rN}^{T})^{T}\in\mathbb{R}^{N\times 3}$ in the rotation directions and $\boldsymbol{J}_{t}=(\boldsymbol{J}_{t1}^{T}\cdots\boldsymbol{J}_{tN}^{T})^{T}\in\mathbb{R}^{N\times 3}$ in the translation directions, where $\boldsymbol{J}_{ri}\in\mathbb{R}^{1\times 3}$ is the Jacobian of the i-th residual $f_{i}\boldsymbol{(}x)$ w.r.t. the 3-DOF rotation and $\boldsymbol{J}_{ti}\in\mathbb{R}^{1\times 3}$ is the Jacobian of the i-th residual $f_{i}\boldsymbol{(}x)$ w.r.t. the 3-DOF translation. That is, it is equivalent to performing eigenvalue decomposition on the top-left 3$\times$3 matrix $\boldsymbol{H}_{r}$ and the bottom-right 3$\times$3 matrix $\boldsymbol{H}_{t}$ of the Hessian matrix $\boldsymbol{H}$, as follows:
@@
$\lambda_{j}=\sum_{i=1}^{N}(\boldsymbol{J}_{ri}\boldsymbol{v}_{rj})^{2}$
@@
$\lambda_{j}=\sum_{i=1}^{N}(\boldsymbol{J}_{ti}\boldsymbol{v}_{tj})^{2}$
@@ IV-B3 Localizability Contribution Vector
$\boldsymbol{F}_{r}=\left[\boldsymbol{F}_{r1}^{e}\cdots\boldsymbol{F}_{rN_{1}}^{e}\;\boldsymbol{F}_{r1}^{p}\cdots\boldsymbol{F}_{rN_{2}}^{p}\right]^{T}$
$\boldsymbol{F}_{t}=\left[\boldsymbol{F}_{t1}^{e}\cdots\boldsymbol{F}_{tN_{1}}^{e}\;\boldsymbol{F}_{t1}^{p}\cdots\boldsymbol{F}_{tN_{2}}^{p}\right]^{T}$
@@ IV-C1 Filtering Out Noise and Picking Out High Localizability Contributions
An example of the calculation results for $L_{f}(j)$ and $L_{u}(j)$ is provided in Fig. 3. They are computed from the localizability contributions of a LiDAR scan performed in an open area with slightly undulating terrain. Fig. 3(a) shows the edge points and planar points extracted from this LiDAR scan. Figs. 3(b) and (c) present the corresponding calculation results. In the directions along $\boldsymbol{v}_{r1}$, $\boldsymbol{v}_{r2}$ and $\boldsymbol{v}_{t1}$, the values of $L_{f}(j)$ and $L_{u}(j)$ are large, indicating well-constrained conditions. In the directions along $\boldsymbol{v}_{r3}$, $\boldsymbol{v}_{t2}$ and $\boldsymbol{v}_{t3}$, the values of $L_{f}(j)$ and $L_{u}(j)$ are small, indicating ill-conditioned constraints. In ill-conditioned directions, the calculation of $L_{u}(j)$ which represents the sum of high localizability contributions, is dominated by the contributions from edge point-to-line correspondences (red region). Therefore, it is essential to utilize the localizability information from edge point-to-line correspondences.
Figure 3: (a) Edge points (green) and planar points (purple) extracted by the front end of LVI-SAM[11]. These points are from a LiDAR scan obtained in an open area with ground undulations. (b)-(c) Two exemplary double Y-axis histograms presenting $L_{f}$ and $L_{u}$ along each eigenvector direction of the 6-DOF pose. The localizability contributions are computed from the valid points shown in (a). It indicates smaller values of $L_{f}(j)$ and $L_{u}(j)$ in the $\boldsymbol{v}_{r3}$, $\boldsymbol{v}_{t2}$ and $\boldsymbol{v}_{t3}$ directions. The red region represents the localizability contribution from edge point-to-line correspondences, while the orange region represents the contribution from planar point-to-plane correspondences.
@@ IV-C2 Categorization
That is, when $L_{f}(j)\geq T_{1}\text{ or }L_{u}(j)\geq T_{2}$, the localizability category of the direction along $\boldsymbol{v}_{j}$ is considered fully localizable, i.e. the category is “Full”. When $L_{f}(j)\geq T_{3}\text{ and }L_{u}(j)\geq T_{4}$, the localizability category of the direction along $\boldsymbol{v}_{j}$ is considered partially localizable, i.e. the category is “Partial”. In this direction, both $L_{f}(j)$ and $L_{u}(j)$ must meet their respective thresholds. This means that there must be a sufficient number of points with high localizability contributions in the direction along $\boldsymbol{v}_{j}$. Otherwise, the localizability category of the direction of $\boldsymbol{v}_{j}$ is considered non-localizable, i.e. the category is “None”. In this study, based on experimental validation, for subsequent experiments, $T_{1}$ is set to 50, $T_{2}$ is set to 30, $T_{3}$ is set to 15 and $T_{4}$ is set to 9.
@@
The categories for directions along $\boldsymbol{v}_{r1}$, $\boldsymbol{v}_{r2}$ and $\boldsymbol{v}_{t1}$ are fully localizable, indicating good geometric constraints. the categories for directions along $\boldsymbol{v}_{r3}$, $\boldsymbol{v}_{t2}$ are partially localizable, suggesting that although these directions exhibit degeneracy, some geometric information may still be usable. The category for directions along $\boldsymbol{v}_{t3}$ is non-localizable, indicating severe degeneracy in this direction with minimal usable geometric information.
Figure 4: An example of the Localizability Detection Module. (a) Scan: A LiDAR scan obtained in an open area with slightly undulating terrain. The same point cloud as shown in Fig. 3 (a) is visualized here from a different viewpoint. (b) Localizability Contribution: Two double Y-axis histograms identical to those in Fig. 3. They are calculated based on the localizability contribution of the scan in (a). (c) Categorization: The localizability category along each eigenvector direction of the 6-DOF pose is assigned through conditional judgment. The threshold is set according to Section IV-C. (d)Result: The results of the localizability categories along the eigenvector directions are shown.
@@ V Optimization Module
The Optimization Module utilizes the localizability information provided by the LocDetect-Module. Based on the localizability categories of the 6-DOF pose along the directions of the eigenvectors, additional soft and hard constraints are incorporated into the optimization equation. Specifically, in partially localizable directions, soft constraints are added using constraint values computed from higher localizability contributions. In non-localizable directions, hard constraints are applied to prevent pose updates. In fully localizable directions, pose updates remain unrestricted.
Note that to improve time efficiency in the subsequent experiments, the localizability category determination and additional constraint calculations are performed only during the first iteration of each ICP execution. For all subsequent iterations before convergence, the categories and additional constraint information computed in the first iteration are used directly in the optimization. Because in most SLAM algorithms, ICP-based pose estimation is typically performed at a relatively high frequency (e.g. 10 Hz in LVI-SAM), this configuration has little impact on accuracy.
@@ V-A Optimization Strategy
Binary degeneracy handling methods may omit usable constraint information, and the use of standalone hard constraints can sometimes induce fluctuations in the trajectory along partially localizable directions. To address the above issue, we use an optimization method that combines soft and hard constraints in the Optimization-Module, utilizing the localizability information from the LocDetect-Module. Each time the ICP algorithm starts, the process begins with the initial pose estimate $\boldsymbol{x}_{0}$, which can be obtained through various methods in SLAM, such as previous odometry. Through iteration, the goal is to find the optimal pose update $\Delta\boldsymbol{x}$, which is added to $\boldsymbol{x}_{0}$ to obtain the optimal pose solution $\boldsymbol{x}^{*}$.
@@
    Partially Localizable (Partial): First, obtain the edge point-to-line correspondences and planar point-to-plane correspondences with moderate localizability contributions in the direction of $\boldsymbol{v}_{j}$ from the LocDetect-Module (i.e., satisfying (27)). Then, use them for a small local ICP optimization, which combines point-to-line and point-to-plane cost functions, to obtain an additional constraint value $\Delta\boldsymbol{r}_{0}\in\mathbb{R}^{3\times 1}$ or $\Delta\boldsymbol{t}_{0}\in\mathbb{R}^{3\times 1}$ for the pose update in the direction of $\boldsymbol{v}_{j}$. Then, a soft constraint is added with weights. The specific process is described in Section V-B.
@@ V-B Adding Soft Constraints in Partially Localizable Directions
    A soft constraint is added to the final optimization equation, encouraging the final pose update in the direction of $\boldsymbol{v}_{j}$ to approach the projection of the constraint value $\Delta\boldsymbol{r}_{0}$ or $\Delta\boldsymbol{t}_{0}$ in the direction of $\boldsymbol{v}_{j}$. A squared loss function is used, as follows:
@@
$\mu_{j}$ is the set weight. By adjusting the value of $\mu_{j}$, the extent to which the final pose update’s projection in the $\boldsymbol{v}_{j}$ direction tends toward the projection of the constraint values $\Delta\boldsymbol{r}_{0}$ or $\Delta\boldsymbol{t}_{0}$ in the $\boldsymbol{v}_{j}$ direction can be controlled. In this paper, $T_{5}$ is set to 15 for all subsequent experiments. When $L_{u}(j)$ in the direction of $\boldsymbol{v}_{j}$ is greater than $T_{5}$, it indicates that more correspondences with high localizability contributions are used to compute $\Delta\boldsymbol{r}_{0}$ or $\Delta\boldsymbol{t}_{0}$, making the calculated values more reliable, and thus the weight $\mu_{j}$ is increased. If the quality and quantity of the correspondences used in (32) or (33) are high, the weight $\mu_{j}$ can be appropriately increased. However, if $\mu_{j}$ is too large, it may cause fluctuations and jaggedness in the optimized trajectory, leading to instability. The setting of $\mu_{j}$ should ensure that the pose estimate in the direction of $\boldsymbol{v}_{j}$ tends toward the constraint value, while also maintaining the smoothness and stability of the trajectory. Nonetheless, the use of soft constraints in partially localizable directions may introduce some uncertainty in the estimation results, and their impact will be further investigated in future work.
@@ VI Experiment
In this section, the experimental setup is introduced in Section VI-A. The proposed method is then evaluated through experiments on a simulated planetary-like environment dataset (Section VI-B), a real-world underground tunnel dataset (Section VI-C) and the Long Corridor sequence from the SubT-MRS dataset (Section VI-D), with comparisons to state-of-the-art methods. Finally, the runtime performance of the algorithm is evaluated (Section VI-E).
@@ VI-A1 Implementation Details
The proposed framework, LP-ICP, is integrated into the point cloud registration framework of LVI-SAM [11]. LVI-SAM combines LIO and VIO sub-modules, with the LIO sub-module serving as the main component, while the VIO sub-module provides initial estimates for the LIO sub-module. The LIO module performs scan-to-map registration using edge point-to-line and planar point-to-plane correspondences. The tested algorithms are run five times on each sequence, and the best result is reported for evaluation. All experiments and evaluations in this paper are performed on a laptop equipped with an Intel i7-12700H CPU.